OpenStudy (anonymous):

State whether the ratios form a proportion. 16 to 18 and 24 to 27 A. Yes B. No

OpenStudy (anonymous):

16/18 = 24/27 ? get to smallest numerator and denominator, each of them 8/9 = 8/9 ? yes